OpenStudy (anonymous):

What are the solution(s) to the following system of equations? y = x2 – 1 y = x – 1

OpenStudy (goformit100):

DO YOU KNOW MATRICES...

OpenStudy (anonymous):

You can just factor the top equation int (x+1)(x-1). Since y=y you know (x+1)(x-1) = (x-1) This equals (x+1)(x-1)-(x-1) = 0 (x-1)x = 0. Thus x = 0 and x = 1 are your two solutions.